Suppose that an economy is characterized by M= $9 trillion V=2 P= base index = 1.0 

Instructions: Enter your responses rounded to two decimal places (do not include a negative sign (-) with your answers). 

a. What is the real value of output (Q)? $D 18 trillion Now assume that the Fed increases the money supply by 20 percent and velocity remains unchanged. 

b. If the price level remains constant, by how much will real output increase? 

c. If, instead, real output is fixed at the natural level of unemployment (= Q from part a), by how much will prices increase in percentage terms? 

d. By how much would V have to decrease to offset the increase in M (assuming Qand P did not change)?
Answer #1

Given,

Money Supply, M = $9 trillion

Velocity, V = 2

P = Base Index = 1.0

According to the Equation of Exchange, If Q is the real output, MV = PQ

a. From the above equation, we obtain Q = MV/P = ($9 trillion x 2)/1.0 = $18 trillion

b. If the Money supply increases by 20%, then M = 1.2 x ($9 trillion) = $10.8 trillion

If M & V remain unchanged, the Real Output = MV/P = ($10.8 trillion x 2)/1.0 = $21.6 trillion

Change in real output = $21.6 trillion - $18 trillion = $3.6 trillion

c. If M = $10.8 trillion, V=2 and Q = $18 trillion

From Equation of Exchange, MV=PQ, we obtain P = MV/Q = ($10.8 trillion x 2)/$18 trillion = 1.2

Percentage increase in prices = (Changed P - Initial P)/Initial P x 100 = (1.2-1.0)/1.0 x 100 = 20%

d. If M = $10.8 trillion, P=1.0 and Q = $18 trillion

From Equation of Exchange, MV=PQ, we obtain, V=PQ/M = (1.0 x $18 trillion)/$10.8 trillion = 1.67

Decrease in V = 2 - 1.67 = 0.33